RESONANT TWO-PHOTON DISSOCIATION SPECTROSCOPY OF FeO^{+}: CHARACTERIZATION OF THE ^{6}Σ GROUND STATE

\underline{RICARDO B. METZ}, Department of Chemistry, University of Massachusetts, Amherst, MA 01003.

We have measured several bands in a ^{6}Π−^{6} Σ transition in FeO^{+} with rotational resolution. In this one-color, two-photon resonant dissociation experiment, one photon excites molecules to the ^{6}Π state, while the second photon dissociates the excited molecules. The spectrum is measured by monitoring the yield of Fe^{+} fragments. Results of the rotational analysis of the (^{6}Π_{7/2}, v=8, 9)-(^{6}Σ, v=0) bands will be presented.